OpenStudy (linstead_cpd_freak):

Solve the word problem. A group of roommates decide to split the cost of a television for their house. If a is the amount spent on the television and n is the number of roommates, the formula C = determines each person’s share of the cost. Each person’s share was $179.60 and 5 roommates live in the house. What was the cost of the television? A. $184.60 B. $679.60 C. $898.00 D. $900.00

OpenStudy (xguardians):

Welcome to OpenStudy, I'd be happy to help you with the question. Would you like to walk through it with me?

OpenStudy (linstead_cpd_freak):

C = a/n

OpenStudy (xguardians):

Price = c * n

OpenStudy (xguardians):

Where?

OpenStudy (xguardians):

@linstead_cpd_freak, use this: Price = c * n Try inputting the values you were given in the problem to solve for the price.

OpenStudy (alones.):

Is this right? This is how I did it.. 179.60 = a/5 a = 179.60 x 5 a = 898

OpenStudy (linstead_cpd_freak):

So $179.60 x 5 = 898.00? Would that make the answer C?

OpenStudy (alones.):

Yes, correct!

OpenStudy (alones.):

You can't medal both of us, so i'll medal @xGuardians for u.

OpenStudy (linstead_cpd_freak):

Thank you

OpenStudy (alones.):

Yep!